# The T2K Collaboration Results from the T2K experiment on neutrino mixing including a new far detector -like sample K. Abe

### v2

@@ Abstract
Acknowledgements—The T2K collaboration would like to thank the J-PARC staff for superb accelerator performance. We thank the CERN NA61/SHINE Collaboration for providing valuable particle production data. We acknowledge the support of MEXT, JSPS KAKENHI and bilateral programs, Japan; NSERC, the NRC, and CFI, Canada; the CEA and CNRS/IN2P3, France; the Deutsche Forschungsgemeinschaft (DFG 397763730, 517206441), Germany; the NKFIH (NKFIH 137812 and TKP2021-NKTA-64), Hungary; the INFN, Italy; the Ministry of Science and Higher Education (2023/WK/04) and the National Science Centre (UMO-2018/30/E/ST2/00441, UMO-2022/46/E/ST2/00336 and UMO-2021/43/D/ST2/01504), Poland; the RSF (RSF 24-12-00271) and the Ministry of Science and Higher Education, Russia; MICINN, ERDF and European Union NextGenerationEU funds and CERCA program and Generalitat de Catalunya, Spain; the SNSF and SERI, Switzerland; the STFC and UKRI, UK; the DOE, USA; and NAFOSTED (103.99-2023.144,IZVSZ2.203433), Vietnam. We also thank CERN for the UA1/NOMAD magnet, DESY for the HERA-B magnet mover system, the BC DRI Group, Prairie DRI Group, ACENET, SciNet, and CalculQuebec consortia in the Digital Research Alliance of Canada, and GridPP in the United Kingdom, the CNRS/IN2P3 Computing Center in France and NERSC, USA. In addition, the participation of individual researchers and institutions has been further supported by funds from the ERC (FP7), “la Caixa” Foundation, the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant; the JSPS, Japan; the Royal Society, UK; French ANR and Sorbonne Université Emergences programmes; the VAST-JSPS (No. QTJP01.02/20-22); and the DOE Early Career programme, USA. For the purposes of open access, the authors have applied a Creative Commons Attribution license to any Author Accepted Manuscript version arising.
@@
P. L. Connolly, S. A. Kahn, M. J. Murtagh, R. B. Palmer, N. P. Samios, and M. Tanaka, Phys. Rev. D 23, 2495 (1981).
@@
T. Kitagaki, M. Etou, H. Sagawa, K. B. McConnel Mahn, E. J. Jeon, and M. Sakuda, in 2nd International Workshop on Neutrino-Nucleus Interactions in the Few GeV
Region (2003).

### v3

@@ Abstract
The T2K collaboration would like to thank the J-PARC staff for superb accelerator performance. We thank the CERN NA61/SHINE Collaboration for providing valuable particle production data. We acknowledge the support of MEXT, JSPS KAKENHI and bilateral programs, Japan; NSERC, the NRC, and CFI, Canada; the CEA and CNRS/IN2P3, France; the Deutsche Forschungsgemeinschaft (DFG 397763730, 517206441), Germany; the NKFIH (NKFIH 137812 and TKP2021-NKTA-64), Hungary; the INFN, Italy; the Ministry of Science and Higher Education (2023/WK/04) and the National Science Centre (UMO-2018/30/E/ST2/00441, UMO-2022/46/E/ST2/00336 and UMO-2021/43/D/ST2/01504), Poland; the RSF (RSF 24-12-00271) and the Ministry of Science and Higher Education, Russia; MICINN (PID2022-136297NB-I00 /AEI/10.13039/501100011033/ FEDER, UE, PID2024-157541NB-I00 (UAM) and PID2023-146401NB-I00 (US), Severo Ochoa Centres of Excellence Programme 2025-2029 (CEX2024001442-S)), Government of Andalucia (FQM160) and the University of Tokyo ICRR’s Inter-University Research Program FY2025 Ref. J1, and ERDF and European Union (UAM: H2020-MSCA-RISE-GA872549- SK2HK) and NextGenerationEU funds (PRTR-C17.I1) and Generalitat de Catalunya (AGAUR 2021-SGR-01506, CERCA program), University of Seville grant (RYC2022-035203-I funded by MICIU/AEI/10.13039/501100011033, “ERDF a way of making Europe” and FSE+, Ayudas “Atracción de Investigadores con Alto Potencial” Ref. VIIPPIT-2025), and Secretariat for Universities and Research of the Ministry of Business and Knowledge of the Government of Catalonia and the European Social Fund (2022FI_B 00336), Spain; the SNSF and SERI, Switzerland; the STFC and UKRI, UK; the DOE, USA; and NAFOSTED (103.99-2023.144,IZVSZ2.203433), Vietnam. We also thank CERN for the UA1/NOMAD magnet, DESY for the HERA-B magnet mover system, the BC DRI Group, Prairie DRI Group, ACENET, SciNet, and CalculQuebec consortia in the Digital Research Alliance of Canada, and GridPP in the United Kingdom, the CNRS/IN2P3 Computing Center in France and NERSC, USA. In addition, the participation of individual researchers and institutions has been further supported by funds from the ERC (FP7), “la Caixa” Foundation, the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant, the Horizon Europe Marie Sklodowska-Curie Staff Exchange project JENNIFER3 grant 101183137; the JSPS, Japan; the Royal Society, UK; French ANR and Sorbonne Université Emergences programmes; the VAST-JSPS (No. QTJP01.02/20-22); and the DOE Early Career program, USA.
Abi et al. [2020]B. Abi et al. (DUNE), Deep Underground Neutrino Experiment (DUNE), Far Detector
Technical Design Report, Volume II: DUNE Physics, arXiv:2002.03005 [hep-ex] (2020).
Brun et al. [1994]R. Brun, F. Bruyant,
F. Carminati, S. Giani, M. Maire, A. McPherson, G. Patrick, and L. Urban, GEANT
Detector Description and Simulation Tool (1994).
Abgrall et al. [2016a]N. Abgrall et al. (NA61/SHINE), Measurements of ,
, , and proton production in
proton–carbon interactions at 31 GeV/c with the NA61/SHINE
spectrometer at the CERN SPS, Eur. Phys. J. C 76, 84 (2016a), arXiv:1510.02703 [hep-ex] .
Abgrall et al. [2016b]N. Abgrall et al. (NA61/SHINE), Measurements of
differential yields from the surface of the T2K replica target for incoming
31 GeV/c protons with the NA61/SHINE spectrometer at the CERN SPS, Eur. Phys. J. C 76, 617 (2016b), arXiv:1603.06774 [hep-ex] .
Abgrall et al. [2019]N. Abgrall et al. (NA61/SHINE), Measurements of , and proton double differential yields from the surface of the T2K replica
target for incoming 31 GeV/c protons with the NA61/SHINE spectrometer at the
CERN SPS, Eur. Phys. J. C 79, 100 (2019), arXiv:1808.04927 [hep-ex]
Ankowski et al. [2015]A. M. Ankowski, O. Benhar, and M. Sakuda, Improving the accuracy of neutrino
energy reconstruction in charged-current quasielastic scattering off nuclear
targets, Phys. Rev. D 91, 033005 (2015).
Chakrani et al. [2024]J. Chakrani et al., Parametrized uncertainties in the spectral function model of neutrino
charged-current quasielastic interactions for oscillation analyses, Phys. Rev. D 109, 072006 (2024), arXiv:2308.01838 [hep-ex] .
Rein and Sehgal [1981]D. Rein and L. M. Sehgal, Neutrino Excitation of
Baryon Resonances and Single Pion Production, Annals Phys. 133, 79 (1981).
Wilkinson et al. [2014]C. Wilkinson, P. Rodrigues, S. Cartwright, L. Thompson, and K. McFarland, Reanalysis of bubble
chamber measurements of muon-neutrino induced single pion production, Phys. Rev. D 90, 112017 (2014), arXiv:1411.4482 [hep-ex] .
Rodrigues et al. [2016]P. Rodrigues, C. Wilkinson, and K. McFarland, Constraining the GENIE
model of neutrino-induced single pion production using reanalyzed bubble
chamber data, Eur. Phys. J. C 76, 474 (2016), arXiv:1601.01888 [hep-ex]
Derrick et al. [1981]M. Derrick et al., Study of
single-pion production by weak neutral currents in low-energy
interactions, Phys. Rev. D 23, 569 (1981).
Radecky et al. [1982]G. M. Radecky et al., Study of
single-pion production by weak charged currents in low-energy
interactions, Phys. Rev. D 25, 1161 (1982).
@@
P. L. Connolly, S. A. Kahn, M. J. Murtagh, R. B. Palmer, N. P. Samios, and M. Tanaka, Study of the isospin structure of single-pion production in
charged-current neutrino interactions, Phys. Rev. D 23, 2495 (1981).
Kitagaki et al. [1986]T. Kitagaki et al., Charged-current exclusive pion production in neutrino-deuterium
interactions, Phys. Rev. D 34, 2554 (1986).
@@
T. Kitagaki, M. Etou, H. Sagawa, K. B. McConnel Mahn, E. J. Jeon, and M. Sakuda, BNL
7-foot bubble chamber experiment: Neutrino deuterium interactions, in 2nd International Workshop on
Neutrino-Nucleus Interactions in the Few GeV Region (2003).
Aguilar-Arevalo et al. [2011a]A. A. Aguilar-Arevalo et al. (MiniBooNE), Measurement of Neutrino-Induced
Charged-Current Charged Pion Production Cross Sections on Mineral Oil at
E, Phys. Rev. D 83, 052007 (2011a), arXiv:1011.3572 [hep-ex] .
Aguilar-Arevalo et al. [2010]A. A. Aguilar-Arevalo et al. (MiniBooNE), Measurement of and
induced neutral current single production cross
sections on mineral oil at , Phys. Rev. D 81, 013005 (2010), arXiv:0911.2063 [hep-ex] .
Aguilar-Arevalo et al. [2011b]A. A. Aguilar-Arevalo et al. (MiniBooNE), Measurement of -induced
charged-current neutral pion production cross sections on mineral oil at
GeV, Phys. Rev. D 83, 052009 (2011b), arXiv:1010.3264 [hep-ex] .
Abe et al. [2020]K. Abe et al. (T2K), Measurement of the muon neutrino charged-current single
production on hydrocarbon using the T2K off-axis near detector
ND280, Phys. Rev. D 101, 012007 (2020), arXiv:1909.03936 [hep-ex] .
Niewczas and Sobczyk [2019]K. Niewczas and J. T. Sobczyk, Nuclear transparency in
monte carlo neutrino event generators, Phys. Rev. C 100, 015505 (2019).
Abe et al. [2017]K. Abe et al. (The T2K Collaboration), Measurement of neutrino and
antineutrino oscillations by the t2k experiment including a new additional
sample of interactions at the far detector, Phys. Rev. D 96, 092006 (2017).
Zyla et al. [2020]P. A. Zyla et al. (Particle Data Group), Review of Particle Physics, PTEP 2020, 083C01 (2020), and 2021
update.
Abe et al. [2014]Y. Abe et al. (Double Chooz), Improved measurements of the neutrino mixing
angle with the Double Chooz detector, JHEP 10, 086, [Erratum: JHEP 02, 074 (2015)], arXiv:1406.7763 [hep-ex] .
Bak et al. [2018]G. Bak et al. (RENO), Measurement of reactor antineutrino oscillation amplitude
and frequency at reno, Phys. Rev. Lett. 121, 201801 (2018), arXiv:1806.00248 [hep-ex] .
Abe et al. [2025a]K. Abe et al. (T2K), Testing T2K’s Bayesian constraints with priors in
alternate parameterisations, arXiv:2507.02101 [hep-ex] (2025a).
Metropolis et al. [1970]N. Metropolis, A. W. Rosenbluth, M. N. Rosenbluth, A. H. Teller, and E. Teller, Equation of state
calculations by fast computing machines, The Journal of Chemical Physics 21, 1087 (1970).
Gelman et al. [1996]A. Gelman, X.-L. Meng, and H. Stern, Posterior predictive assessment of model fitness
via realized discrepancies, Statistica sinica , 733 (1996).
Gelman et al. [2000]A. Gelman, Y. Goegebeur,
F. Tuerlinckx, and I. V. Mechelen, Diagnostic checks for discrete data
regression models using posterior predictive simulations, Appl. Statist. 49 Par 2, 247 (2000).
Martini et al. [2009] M. Martini, M. Ericson, G. Chanfray, and J. Marteau, A
Unified approach for nucleon knock-out, coherent and incoherent pion
production in neutrino interactions with nuclei, Phys. Rev. C 80, 065501 (2009), arXiv:0910.2622 [nucl-th] .
Jachowicz et al. [2002]N. Jachowicz, K. Heyde,
J. Ryckebusch, and S. Rombouts, Continuum random phase approximation approach to
charged current neutrino nucleus scattering, Phys. Rev. C 65, 025501 (2002).
Pandey et al. [2015]V. Pandey, N. Jachowicz,
T. Van Cuyck, J. Ryckebusch, and M. Martini, Low-energy excitations and quasielastic contribution to
electron-nucleus and neutrino-nucleus scattering in the continuum
random-phase approximation, Phys. Rev. C 92, 024606 (2015), arXiv:1412.4624 [nucl-th] .
Jarlskog [1985]C. Jarlskog, Commutator of the quark
mass matrices in the standard electroweak model and a measure of maximal
nonconservation, Phys. Rev. Lett. 55, 1039 (1985).
Krastev and Petcov [1988]P. I. Krastev and S. T. Petcov, Resonance Amplification
and t Violation Effects in Three Neutrino Oscillations in the Earth, Phys. Lett. B 205, 84 (1988).
Abe et al. [2025b]K. Abe et al. (The T2K Collaboration), Markov chain monte carlo chain release
for "results from the t2k experiment on neutrino mixing including a new far
detector muon-like sample", 10.5281/zenodo.17312780 (2025b).